A runner finishes one lap if the distance from starting point to finish line is 50 m and he finishes the love within one minute what is his average velocity
1
Expert's answer
2021-09-23T07:58:13-0400
"x_1=0;x_2=50m;t_1=0sec;t_2=60sec"

We know that

Average velocity

"V_{ave}=\frac{x_2-x_1}{t_2-t_1}"

"V_{ave}=\frac{50-0}{60-0}=0.833m/sec"